Can someone please help me solve the quadratic equation
7x^2-5=0
This is what I have tried, but I don't think I am on the right track so I didn't go any further.
x(7x+5)=0
x=0 7x+5=0
7x=-5
x=-5/7

Answer by bobdabuildr(27)

7x^2-5=0
7x^2=5
x^2=5/7
x=sqrt(5/7)